# Efficient Multimodal Learning from Data-centric Perspective

### v2

@@ 4.3 Ablation Study
We also investigate Bunny’s performance under different choices of vision encoder and language model. As shown in Tab. 5, all combinations show exceptional performance, with SigLIP-SO [14] and Phi-3 [13] delivering the best results. This indicates the customizability of the Bunny framework and the high quality of the constructed training data.
@@ 5 Conclusion
We present Bunny, a family of lightweight but powerful multimodal models. It offers multiple plug-and-play vision encoders, including EVA-CLIP and SigLIP, and language backbones, including Phi-1.5, Qwen1.5-1.8B, StableLM-2, MiniCPM-2B, Phi-2, Phi-3-Mini, and Llama-3-8B. To compensate for the decrease in model size, we construct more informative training data by condensing data from a broader source. Remarkably, our Bunny-4B/8B outperforms the state-of-the-art large MLLMs on multiple benchmarks.

### v3

@@ 4.3 Ablation Study
We also investigate Bunny’s performance with various combinations of vision encoders and language models. As shown in Tab. 12, all combinations show exceptional performance, with SigLIP-SO [14] and Llama-3-8B [21] achieving the best results. This highlights the flexibility of the Bunny framework and the high quality of the curated training data. Note that in the fine-tuning stage, Phi-1.5, StableLM-2 and Phi-2-based Bunny are trained on Bunny-695K for 1 epoch with the vision encoder frozen. In contrast, Phi-3-Mini and Llama-3-8B-based Bunny are trained using additional fine-tuning data, S2-Wrapper, trainable vision encoder, weight merging, etc., as illustrated before.
@@ 5 Conclusion
We present Bunny, a family of lightweight but powerful multimodal models. It offers multiple plug-and-play vision encoders, including EVA-CLIP and SigLIP, and language backbones, including Phi-1.5, Qwen1.5-1.8B, StableLM-2, MiniCPM-2B, Phi-2, Phi-3-Mini, and Llama-3-8B. To compensate for the decrease in model size, we construct more informative training data by selecting data from a broader source. Remarkably, our Bunny-4B/8B outperforms the state-of-the-art large MLLMs on multiple benchmarks.